OpenStudy (anonymous):

use the matrix tool to solve the system of equations. 8x+5y=9 -x+y=7

OpenStudy (jdoe0001):

so have you covered "inverse matrices" yet?

OpenStudy (anonymous):

what's that.?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=PTN8rganLWo

OpenStudy (anonymous):

yea I did.

OpenStudy (jdoe0001):

so, what would you get for the determinant for the 2x2 => \(\bf \left[\begin{matrix} 8 & 5 \\ -1 & 1 \end{matrix}\right] \)

OpenStudy (anonymous):

1/(8)(1) 1/(-1)(1) .?

OpenStudy (jdoe0001):

the video covers also how to get determinants :)

OpenStudy (anonymous):

1/ ad-bc

OpenStudy (jdoe0001):

well, that'd be the inverse of it, yes

OpenStudy (jdoe0001):

which you'd be needing too

OpenStudy (jdoe0001):

let us first find the determinant only :)

OpenStudy (anonymous):

kay. then umm..?

OpenStudy (jdoe0001):

so, what did you get for the determinant of \(\bf \left[\begin{matrix} 8 & 5 \\ -1 & 1 \end{matrix}\right]\)

OpenStudy (anonymous):

9&7.?

OpenStudy (jdoe0001):

the video is worth the while, it covers using the inverse determinant to solve the system, and matrix multiplication too :)

OpenStudy (anonymous):

okay. then.

OpenStudy (anonymous):

but when does he talk about the determinant.?

OpenStudy (anonymous):

7-4=3.?